# System 0/1/2/3: Quad-process theory for multi-timescale embodied collective cognitive systems

### v1

@@ 4.4 Timescales of Canalization (System 0/1/2/3)
The concept of canalization encompasses a variety of landscapes, ranging from physics to language. Physical laws can be regarded as an extreme case of canalized landscapes. Physical matter, at least as it was understood in Bergson’s time, does not make ““decisions””; all behavior is predetermined and immutable. This constitutes the physical landscape. Next, on the timescale of biological evolution, canalization defines species-specific landscapes that correspond to their Umwelt. Movements such as passive walking, which are automatically generated by morphological features (System 0), are positioned here. On the timescale of an individual”s lifetime, "learning" further diversifies the folds of the landscape. The mastery of a skill leads to varying degrees of ease in performing actions and manipulating conceptual tools (Systems 1 and 2).
@@
Finally, in humans who engage in social activities using language, canalization also occurs on the timescale of cultural evolution. In his fourth and final major book, The Two Sources of Morality and Religion, Bergson developed the concept of a socially canalized landscape. When people act collectively, certain behavioral channels are carved into the social landscape to resolve coordination problems, such as social norms, rules, and morality. Similar to evolutionary and learning landscapes, this canalization promotes a specific repertoire of behaviors unique to that community. Without this, language use would not be possible (System 3; see Section 2.3).
@@
Common to these different levels of landscape is that, while they act as constraints by reducing possibilities, they also facilitate and promote stereotyped action, reducing or even eliminating the need for the present window. Nevertheless, having a present window and access to unorganized pure memory hold significant importance for understanding the creativity of life.
@@
First, Bergson reflected on the historical process by which the present window was gradually acquired through evolution. In simple organisms, reactions "can then hardly be delayed" ([11], p. 22). Distant perception emerges when organisms can "defer the date of their fulfillment" ([11], p. 23) of responses. The present window and reference to pure memory allow non-random improvisations in unforeseen situations, enabling complex adaptive behaviors that are not available through evolution or learning alone.
@@
Besides, on-the-fly variability promotes learning efficiency. Bergson notes: "thanks to this faculty, we have no need to await at the hands of chance the accidental repetition of the same situations, in order to organize into a habit concomitant movements; we make use of the fugitive image to construct a stable mechanism which takes its place" ([11], p. 98). Similarly, machine learning exhibits improved efficiency with additional working memory ([39], [147]).
@@
While the present window enriches phenomenological time perception, its biological size is limited to a few seconds. The invention of language can be seen as artificially extending this limit, which we call the Extended Window of the Present (EWP). In a later pape, Bergson wrote the following:
@@
"Our consciousness tells us that when we speak of our present we are thinking of a certain interval of duration. What duration? It is impossible to fix it exactly, as it is something rather elusive. My present, at this moment, is the sentence I am pronouncing. But […] my attention is something that can be made longer or shorter, like the interval between the two points of a compass." ([12], pp. 178-179).
@@
Language stands out for its remarkable ability to enhance the resolution and plasticity of human perception and memory, achieving a level of complexity unimaginable for organisms constrained by biological or personal limitations. As Bergson observed in Creative Evolution (1907), humans, unlike socially organized yet instinct-driven hymenopterans, have attained an ““unlimited”” complexity. The present window is extended through various means, such as conversation, symbolic note-taking, transactive memory, and the incorporation of others” cognitive resources during deliberation and dialogue, which, in essence, utilizes interactions with the social and collective landscape as a form of shared working memory, as shown by the CPC hypothesis (see Section 3.2). Thus, the inherently social nature of language is deeply rooted in the dynamic interplay between the landscape and present window.

### v2

@@ 4.4 Timescales of Canalization (System 0/1/2/3)
The concept of canalization encompasses a variety of landscapes, ranging from physical laws to language, and naturally aligns with the hierarchical structure of Systems 0, 1, 2, and 3 proposed in this paper. Physical laws can be regarded as an extreme form of canalized landscapes. Physical matter, at least as understood in Bergson’s era, does not make “decisions”; its behavior is predetermined and immutable. This constitutes the physical landscape. Next, on the timescale of biological evolution, canalization shapes species-specific landscapes that correspond to their respective Umwelt. Movements such as passive walking, which are automatically generated by morphological features (System 0), are positioned here. On the timescale of an individual’s lifetime, learning further diversifies the folds of the landscape. The mastery of a skill leads to varying degrees of ease in performing actions and manipulating conceptual tools (Systems 1 and 2).
@@
Finally, in humans who engage in social activities through symbols, canalization also occurs on the timescale of cultural evolution. In his fourth and final major work, The Two Sources of Morality and Religion, Bergson developed the concept of a socially canalized landscape. When people act collectively, certain behavioral channels are carved into the social landscape to resolve coordination problems, such as social norms, rules, and morality. Similar to evolutionary and individual learning landscapes, societal canalization fosters community-specific repertoires of behavior. Without this level of canalization, language usage would not be possible (System 3; see Section 2.3).
@@
Common to all these levels of landscapes is their dual nature: while they constrain possibilities by reducing choices, they also facilitate and promote stereotyped actions, reducing or even eliminating the need for the present window. Nevertheless, having a present window and access to unorganized pure memory holds significant importance for understanding the creativity inherent in life.
@@
First, Bergson reflects on the historical process by which the present window was gradually acquired through evolution. In simpler organisms, reactions “can then hardly be delayed” ([11], p. 22). Distant perception emerges when organisms become able to “defer the date of their fulfillment” ([11], p. 23) of responses. Thus, the presence of a present window and access to pure memory enables non-random improvisations in unforeseen situations, facilitating complex adaptive behaviors beyond what evolution or individual learning alone can achieve.
@@
Moreover, on-the-fly variability promotes learning efficiency. Bergson notes: “thanks to this faculty, we have no need to await at the hands of chance the accidental repetition of the same situations, in order to organize into a habit concomitant movements; we make use of the fugitive image to construct a stable mechanism which takes its place” ([11], p. 98). Similarly, machine learning exhibits improved efficiency when supplemented by working memory  [38, 146].
@@
While the present window enriches phenomenological perception of time, its biological size is limited to just a few seconds. The invention of symbolic systems can be seen as artificially extending this limit, resulting in what we term the Extended Window of the Present (EWP). In a later paper, Bergson wrote the following:
@@
“Our consciousness tells us that when we speak of our present we are thinking of a certain interval of duration. What duration? It is impossible to fix it exactly, as it is something rather elusive. My present, at this moment, is the sentence I am pronouncing. But [$\ldots$] my attention is something that can be made longer or shorter, like the interval between the two points of a compass.” ([12], pp. 178-179).
@@
Language stands out for its remarkable capacity to enhance the resolution and plasticity of human perception and memory, enabling a level of complexity unimaginable for organisms constrained by biological or personal limits. As Bergson observed in Creative Evolution (1907), humans, unlike the socially organized yet instinct-driven hymenopterans, have reached an “unlimited” complexity. The present window can be extended through various means, such as conversation, symbolic note-taking, transactive memory, and incorporating cognitive resources of others during deliberation and dialogue. Such interactions with the social and collective landscape effectively serves as shared working memory, as illustrated by the CPC hypothesis (see Section 3.2). Thus, the inherently social nature of language is deeply rooted in the dynamic interplay between the landscape and the present window.